F(x)= Arc cotx² + Arc sinx² . x²
1
Expert's answer
2020-04-08T17:21:47-0400

"F(x)=Arccot(x^2)+x^2Arcsin(x^2)"

"\frac{dF}{dx}=-\frac{1}{1+(x^2)^2}\frac{d}{dx}(x^2)+2xArcsin(x^2)+\frac{x^2}{\sqrt{1-(x^2)^2}}\frac{d}{dx}(x^2)="

"=-\frac{2x}{1+(x^2)^2}+2xArcsin(x^2)+\frac{2x^3}{\sqrt{1-(x^2)^2}}"